# Skill Discovery for Software Scripting Automation via Offline Simulations with LLMs

### v1

@@ 7 Conclusions
In this work, we proposed an offline simulation framework for skill discovery in software scripting automation using LLMs. Unlike prior approaches that rely on runtime code generation, our method pre-build a diverse and verified skillset in two phases: (1) task creation, guided by both top-down functionality categories and bottom-up API synergy exploration, and (2) skill generation with trials, where LLM-generated scripts are iteratively refined based on execution feedback.
Our experiments using Adobe Illustrator as a testbed demonstrated the effectiveness of this approach, achieving a significantly higher success rate while largely reducing response time and token cost by avoiding runtime generation. The top-down task creation strategy encourages broad functional coverage, while the bottom-up API synergy modeling explores the of discovered skills, particularly for underutilized API endpoints. Furthermore, our GCN-based API link prediction model outperformed naive semantic matching in identifying synergistic API pairs, enhancing task creation during the bottom-up simulation.
To the best of our knowledge, this is the first attempt to use software scripting interfaces as a testbed for LLM-based systems, highlighting the advantage of obtaining direct execution feedback in a controlled environment. We identified several future directions, including enhancing LLMs’ knowledge of specific software, leveraging real-world usage data, and generalizing the approach to other API orchestration scenarios. We hope that this work not only provides insights into improving software scripting interfaces but also contributes to bridging the gap between AI capabilities and user needs in specialized domains.
@@
Despite the promising results of our offline simulation framework for skill discovery in software scripting automation, there are several limitations to consider beyond those discussed in Section 6.
First, our method relies heavily on publicly available scripting guides and the existing knowledge of LLMs about the software. In cases where such documentation is incomplete, outdated, or unavailable, or when LLMs have limited knowledge of the software, the effectiveness of our approach may be reduced. This dependence may restrict the applicability of our framework to well-documented software systems and those familiar to the LLMs.
Second, our evaluation primarily focuses on automated success rates and performance metrics within a controlled environment. While these metrics provide valuable insights into technical performance, they do not capture user experience or satisfaction. Future studies could include user evaluations to assess the usability and usefulness of the generated skills with real users.
Third, we expect that queries from real users may be more complex and not always match the granularity of the skills in our skillset. Real-world user queries might be ambiguous or involve higher-level tasks that require combining multiple skills or adapting existing ones. The skillset generated by our framework does not account for individual user preferences or specific workflow requirements. As a result, while our skillset covers a broad range of tasks, it may not fully address the unique needs of all users, limiting its adaptability and personalization.
@@ Appendix A Adobe Illustrator Functionalities
Part of the functionality categories in Illustrator are shown in Table 8.
Category Subcategories
Drawing Drawing basics, Draw pixel-perfect art, Edit paths, …
Color Adjust Colors, Select Colors, Use the Adobe Color Themes panel, …
Painting Gradients, Paint with fills and strokes, Brushes, …
Select and arrange objects Select objects, Move and align objects, Layers, …
Reshape objects Crop images, Transform objects, Puppet Warp, …
Import, export, and save Save artwork, Export artwork, Package files, …
Type Create text, Using fonts in Illustrator, Format paragraphs, …
Create special effects Work with effects, Graphic styles, Drop shadows, …
Web graphics Best practices, Create animations, SVG, …
Printing Set up documents, Print with color management, Overprint, …
Table 8: Illustrator Functionality Overview
@@ Appendix B Prompts
In this section, we show system prompts and user prompts for task creation, skill generation with trials, and LVLM-based validation in Table 9, Table 10, and Table 11, respectively.
System Prompt for Task Creation in Illustrator You are an expert user for Adobe Illustrator. Your goal is to generate as many tasks as possible that are helpful and represent common needs for Illustrator users. The generated tasks should follow the following criteria: 1. Describe these tasks so that they can be coded into a script. 2. The tasks should not be already implemented in Illustrator. 3. The tasks should be minimally dependent on the content. Generate the tasks in plain text. Each task takes one line. Do no generate any other information such as numbering.
User Prompt for Task Creation with Top-down Simulation Give me 10 most useful tasks related to {subcategory} under the category of {category}, in Adobe Illustrator. Examples of successful tasks in the previous rounds include: {a list of successful task descriptions from previous round under the same category}
User Prompt for Task Creation with Bottom-up Simulation Give me 10 most useful Adobe Illustrator tasks related to {api} whose description is: {Names and descriptions of top k synergistic APIs} Take inspiration from the following APIs and their descriptions by considering the possibility of using {api} with at least one of the following APIs. It’s okay to not use them as long as the tasks related to {api} are useful. {top_nodes_info} The generated tasks should follow the following guidelines: - Make sure the tasks are reusable. - The tasks should be logical and reasonable to use two or more APIs together. - The generated task should not simply be a concatenation of two API nodes, i.e., do task A and do a separate task B that doesn’t closely depend on task A. - Prioritize the usefulness of the tasks over generating exactly 10 tasks—fewer, high-quality tasks are acceptable.
Table 9: Prompts for Task Creation in Illustrator
System Prompt for Code Generation You are an assistant generating ExtendScript code for Adobe Illustrator. You will be provided a query that attempts to perform an action in Illustrator. Return only the ExtendScript code snippet without additional messages, formatting, or markdown. Initialize a document to simulate this code. Generate the initialization code and task code separately in the following JSON format: {"init_code": INITIALIZATION_CODE, "code": CODE, "code_name": "[brief description]"} The code must follow these rules: 1. Do not use alert. Return messages for stdout. 2. If the task is not feasible, return {"code": ""}. 3. Start error messages with "Error: ". 4. Include necessary initialization for selecting objects. 5. Ensure reusability of task code. 6. Call the function you create to execute the task. 7. Keep the initial layout minimal for clear visual results. 8. Do not crash Illustrator.
Table 10: System Prompt for ExtendScript Code Generation
System Prompt for Skill Validation You will be given a task description, a piece of initialization code, a layout after running the initialization code, a piece of task code, and a layout after running the task code. The context for the task is Adobe Illustrator. Your job is to judge whether the task was performed correctly or not, given the task description and the two layout figures. The difference between the two layout figures should reflect the result of running the task code. Sometimes, the failure reason can be that the initialization code does not generate necessary elements for the task code to run correctly, or the task code does not perform the task correctly. The output should be in JSON format: {"valid": true/false, "reason": [brief reason for the judgment], "suggestion": [brief suggestion for improvement]} Ensure the output contains no additional messages, formatting, or markdown, so that it can be directly parsed by json.loads().
User Prompt for Validation with LVLM Task description: {task_description} Initialization code: {init_code} Task code: {code} {init_layout.png} {outcome_layout.png}
Table 11: Prompts for Skill Validation

### v2

@@ 6 Conclusions
We proposed an offline simulation framework for skill discovery in software scripting automation with LLMs. Our method pre-build a diverse and verified skillset in two phases: (1) task creation, guided by both top-down functionality categories and bottom-up API synergy exploration, and (2) skill generation with trials, where LLM-generated scripts are iteratively refined through feedback based on execution error and functional outcome.
Experiments on Adobe Illustrator demonstrated that our method achieves higher success rates, broader functional coverage, and lower response time compared to runtime generation. The top-down task creation strategy encourages broad functional coverage, while the bottom-up API synergy modeling explores the of discovered skills, particularly for underutilized API endpoints. Furthermore, our GCN-based API link prediction model outperformed naive semantic matching in identifying synergistic API pairs, enhancing task creation during the bottom-up simulation.
To our best knowledge, this is the first attempt to use software scripting interfaces as a testbed for LLM-based systems, highlighting the advantage of obtaining direct execution feedback in a controlled environment and providing valuable insights into the design of more capable and user-aligned scripting systems.
@@
Despite the promising results of our offline simulation framework for skill discovery in software scripting automation, there are several limitations.
First, our method relies heavily on publicly available scripting guides and the existing knowledge of LLMs about the software. In cases where such documentation is incomplete, outdated, or unavailable, or when LLMs have limited knowledge of the software, LLMs may hallucinate API usage or rely on superficial cues Patil et al. (2024); Zhou et al. (2024); Liu et al. (2025), reducing effectiveness and robustness. This dependence restricts the applicability of our framework to well-documented software systems and those familiar to the LLMs. To address this, one promising direction is to train a domain-specific LLM tailored to the software through iterative trial and error, which could generate more accurate and comprehensive skills. Such methods are complementary to our framework.
Second, our evaluation primarily focuses on automated success rates and performance metrics within a controlled environment. While these metrics provide valuable insights into technical performance, they do not capture user experience or satisfaction. Automatic evaluation can be less reliable on such subjective or high-inference criteria Liu et al. (2016); Amershi et al. (2019); Xu et al. (2024). Although we ran a small-scale user study to evaluate the perceived usefulness of the explored tasks, future studies could include more comprehensive user evaluations to assess the usability and overall impact of the generated skills in real-world workflows. Additionally, mining API interactions from actual scripting projects could provide valuable insights into common user behaviors and task patterns, making the generated skills more applicable in practice. However, obtaining such data is challenging and raises interesting questions about what kinds of behavioral data are useful and how to collect them responsibly.
Third, we expect that queries from real users may be more complex and not always match the granularity of the skills in our skillset. Real-world user queries might be ambiguous or involve higher-level tasks that require combining multiple skills or adapting existing ones. The skillset generated by our framework does not account for individual user preferences or specific workflow requirements. However, the proposed RAG variant shows promising results, significantly improving the success rate with only a modest increase in response time, suggesting its potential to better handle such complex or personalized queries.
Fourth, due to limited support for external programmatic control, we are only able to conduct a large-scale study on Adobe Illustrator. Although we conducted initial exploration on Excel, future work should extend the framework to other scripting environments with broader automation support to validate its generalizability.
Acknowledgments
We would like to thank our collaborators at Adobe Research and the University of Maryland for their helpful discussion, including Dang Minh Nguyen, Dayeon Ki, Nishant Deepa Balepur, Vishakh Padmakumar, Yoonjoo Lee, and Hyunji Lee.
@@ Appendix A Adobe Illustrator Functionalities
Part of the functionality categories in Illustrator are shown in Table LABEL:tab:illustrator_features.
@@ Appendix B Prompts
In this section, we show system prompts and user prompts for task creation, skill generation with trials, and LVLM-based validation in Table LABEL:prompt:creation, Table LABEL:prompt:sys_code, and Table LABEL:prompt:validation, respectively.
Appendix C Example Illustrator Skills
We show an example skill generated by the top-down strategy in Table 12, and one by the bottom-up strategy in Table LABEL:tab:full_example_2.
Appendix D Annotation for Perceived Helpfulness
We asked the two annotators to rate the usefulness of 150 sampled tasks on a 3-point scale: 1 (very rarely useful), 2 (occasionally helpful), and 3 (very useful). Annotators judged each task based solely on its description, and the tasks are randomly shuffled without disclosing We adopted a stricter interpretation of the scale, acknowledging that most tasks offer some utility in niche scenarios, making it difficult to classify any task as completely useless.
The annotators achieved substantial agreement on their ratings, with similar average scores across conditions and only minor differences between them, as shown in Table 9. While exact agreement occurred in $41\%$ of cases, the one-off agreement was remarkably high at $95\%$, indicating that when disagreements occurred, they typically differed by only one point on the scale. This is further supported by the weighted Cohen’s Kappa score of $0.176$, which accounts for the ordinal nature of our rating scale and the fact that most disagreements were minor. Krippendorff’s Alpha (0.171) similarly reflects this pattern of agreement. Overall, these metrics demonstrate that despite the inherent subjectivity in usefulness judgments, our annotators maintained reasonable consistency in their evaluations.
Appendix E Exploring Excel as the Testbed
Test Set Top-down Bottom-up
Rater 1 2.42 2.40 1.78
Rater 2 2.53 2.16 1.72
Table 9: Average usefulness scores (1–3 scale) of sampled tasks, as rated by human evaluators.
Task # Trials
Combine multiple Excel tables into one 5
Count blank rows on all sheets Failed
Return table data as JSON Failed
Remove hyperlinks from each cell 5
Move rows using range values 1*
Table 10: Number of LLM code generation trials required to successfully complete each Excel scripting task. “Failed” indicates that the LLM was unable to produce a working script for the task within five attempts. * due to the ambiguity in task description, it didn’t reproduce the example code in the tutorial but was able to achieve something reasonable.
Model Strategy %Success@1 %Success@3
Llama-3.1-70B Top-down 5.5% 16.8%
Llama-3.1-70B Bottom-up 8.7% 23.5%
deepseek-r1 Top-down 6.9% 27.6%
deepseek-r1 Bottom-up 8.7% 34.8%
Table 11: Success rate at the first and third trials for each model and strategy.
While Adobe Illustrator provides a flexible scripting interface that allows external automation of tasks, Excel’s scripting support is primarily designed for use within its own environment and does not readily permit programmatic control from outside applications. As a result, the approach used to automate and evaluate scripts in Illustrator could not be directly applied to Excel. This limitation restricted our ability to conduct comparable experiments across both platforms.
However, we explore the following aspects to generalize our framework to Excel. (1) API analysis. We collected 2,140 APIs and identified 286 sample scripts from the official tutorial,  [^5] https://learn.microsoft.com/en-us/office/dev/scripts/overview/excel which cover 124 unique APIs. This supports our broader observation that modern software exposes extensive scripting APIs, while available example scripts typically cover only a small subset. (2) LLM-based code generation. We selected five tasks from the “quick scenarios” category in the Excel scripting tutorial and prompted LLM to generate scripts following our framework (see Section 3.2). We show the system prompt for Excel in Table LABEL:prompt:sys_code_excel and user prompt follows the one in Table LABEL:prompt:sys_code. These tasks are aligned with the complexity and skillset expected in our Illustrator experiments. Table 10 reports the number of LLM code generation trials required for each task. We observed that some tasks could be completed with only a few LLM prompt iterations, while others consistently failed, likely due to differences in task difficulty and model familiarity with Excel’s scripting environment. Importantly, two out of five tasks require multiple trials highlights the effectiveness of our offline simulation framework and the trial-and-error approach. Overall, these results indicate both the challenges and opportunities in generalizing LLM-based code generation across software platforms with distinct automation interfaces, and demonstrate that our framework can be extended beyond its initial domain.
Desc arrange selected objects in circle
# Trial No. 2
[…]
Effect [Uncaptioned image] $\Rightarrow$ [Uncaptioned image]
Table 12: Complete version of Table 1. Desc is the natural language description. Effect shows the layout before and after running the skill. Judge is the validation judgment of the LVLM. Code simplified for space constraints.
Desc Alert the user when the name of a tag in the selection matches a predefined list of names.
# Trial No. 2
Code last round (w/ Init. code)